# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 1 Introduction
Figure 3: Taxonomy of LM-based Dialogue System, with representative LMs of each task.
2 Background
2.1 Types of Dialogues
Before we proceed, it is essential to figure out how many different types of dialogs are in our daily life. Table I demonstrates some examples of four major types of dialogs, including:
  1. (a)
    social chit-chat. This type of dialog involves casual and informal conversations aimed at establishing and maintaining social connections. It includes light-hearted discussions about topics such as current events [121, 122], personal experiences [79, 123], hobbies, and interests [124]. Social chit-chat is typically used to create a friendly atmosphere and build social bonds, requiring the DS to generate persona-consistent, empathetic, and compassionate responses besides helpful.
  2. (b)
    knowledge-grounded chit-chat. Knowledge-grounded dialogs revolve around exchanging information and discussing topics that require a certain level of expertise or specific knowledge. It can involve discussing academic subjects, professional domains, or specialized areas of interest. This type of dialog often focuses on sharing and acquiring knowledge [105, 125]. The knowledge can be formed in different formats such as graphs [126], tables [127], figures [128], or even a combination of them [129].
  3. (c)
    question answering. Question-answering dialogs involve a user asking specific questions, and the dialogue system providing relevant and accurate answers. These dialogs are typically focused on obtaining specific information or clarifications. The system’s role is to understand the user’s queries and provide appropriate responses based on the available knowledge or resources [122, 130].
  4. (d)
    task-oriented dialog. Task-oriented dialogs center around assisting users in accomplishing specific tasks or goals. Users seek help from the dialogue system to perform actions such as making reservations, booking flights, ordering products, or seeking recommendations. These dialogs often involve a series of steps or sub-tasks that need to be completed to achieve the overall objective [21, 131, 132].
These are general descriptions of the different types of dialogs, and there may be variations or combinations of these types in real-world dialog scenarios [8]. The diversity and complexity of each type of dialog not only contribute to the wide range of applications for dialogue systems but also give rise to various challenges and problems. These challenges pose unique requirements and considerations, necessitating the development of dialogue systems that are both versatile and robust that can effectively handle the intricacies of each dialog scenario in a unified manner.
2.2 Task Definition
Following the unified definition recently [8, 133], we formulate the task of dialogue system as a seq2seq problem by directly mapping the dialogue context to the final response with or without the assistance of different external sources such as the local database in TOD, different knowledge bases in ODD, e.g., persona and documents. Formally, given the dialogue context $c=\{u_{1},s_{1},...,u_{t}\}$, and optionally the external knowledge bases $K$, the goal of the dialogue system is to generate a helpful and harmless response $r$ as shown below:
$r=\boldsymbol{M}(c,K)\quad or\quad\boldsymbol{M}(c)$ (1)
Where $\boldsymbol{M}$ is parameterized by different language models. Specifically, there are significant differences in interactions between various dialogue systems with the external knowledge base $K$. In this work, we discuss two major interactions corresponding to TOD and ODD respectively.
Formally, the TOD first generates the belief state according to the given context [134, 64]:
$b=\boldsymbol{M}(c)$ (2)
$TOD(r|c,K)=f(r|c,b,s)f(b|c)$ (3)
Secondly, for the ODD, the processing is relatively simpler by directly rewriting the dialogue context as the query or using the whole dialogue context to search the external knowledge base $K$ [105]. In this case, TOD can be regarded as one special type of ODD with the source of knowledge as corresponding database [135].
$ODD(r|c,K)=f(r|c,q,s)f(q|c)$ (4)
It is important to clarify here that ODD contains the cases which does not require any external knowledge, by directly mapping dialogue context to corresponding responses using LMs [136].
3 The Independent Development of TOD & ODD
3.1 The Development of TOD
@@
In the nascent stages of task-oriented dialogue system development, exemplified by systems like GUS [15] and ALICE [16], the technology relied heavily on rule-based approaches. These early systems operated within single domains and adhered to pre-defined decision trees or rule sets. For example, GUS was constrained to specific domains such as flight information retrieval [15], while ALICE was designed to simulate natural language conversation, following a script-like structure based on pattern matching and predefined responses [16]. These systems showcase impressive capabilities within their limited contexts but struggle to handle nuanced or complex conversations. They often require extensive manual crafting of rules and lack the ability to generalize beyond their narrow domains, making them ill-equipped for real-world applications with varying user intents and open-ended dialogues.
@@
With the advent of the neural language model (NLM) and subsequent advancements in recurrent neural networks (RNNs), the landscape of task-oriented dialogue systems underwent a first evolution. The introduction of RNNs allows for more flexible and adaptive responses, enabling systems to process and generate natural language in a contextually aware manner. Dialogue systems now exhibit a broader spectrum of capabilities, embracing multi-domain functionality and accommodating diverse user intents. Unlike their predecessors, these modern systems can handle complex, intertwined tasks within a single conversation. For instance, a task-oriented dialogue system might seamlessly assist a user in booking a flight, reserving a hotel room, and planning an itinerary, all within the same interaction. This expanded capability has been made possible by the integration of four different components: Natural Language Understanding (NLU) [137], Dialogue State Tracking (DST) [138], Dialogue Policy Learning (DPL) [139], and Natural Language Generation (NLG) [140, 141]  [^5] https://github.com/yizhen20133868/Awesome-TOD-NLG-Survey.git.
  • [leftmargin=*,topsep=0pt,itemsep=0pt,parsep=0pt]
  • Natural Language Understanding (NLU): NLU processes user utterances and extracts relevant semantic frames, consisting of two sub-tasks: intent classification and slot filling. NLU can be rule-based, where predefined patterns or rules are used to extract information, or it can employ deep learning techniques like intent classification and named entity recognition.
  • Dialogue State Tracking (DST): DST maintains and updates the state of the conversation as it progresses. It keeps track of the user’s goals, preferences, and any relevant context obtained from the dialogue context. DST can use a rule-based approach to update according to the results from the NLU module, or replace the NLU module and directly track the dialogue state from a dialogue context with similar architecture as NLU [142, 60].
  • Dialogue Policy Learning (DPL): The dialogue policy component determines the system’s actions or responses based on the current dialogue state. It maps the dialogue state to appropriate system actions, which can include asking clarification questions, providing recommendations, or executing relevant API calls. Dialogue policies can be rule-based, handcrafted, or learned using techniques like reinforcement learning [21, 69, 67, 143, 37] or supervised learning [144, 75]. More details can be found in the survey paper [139].
  • Natural Language Generation (NLG): NLG generates system responses in natural language based on the dialogue policy’s output. It transforms structured information or system actions into coherent and fluent sentences that can be understood by users. NLG can employ template-based approaches, rule-based generation, or more advanced techniques like neural network-based generation models [145, 146].
It is worth noting that most studies at this stage independently target different components and propose task-specific designs. Additionally, there are some works try to combine consecutive components, such as word-level DST (coupling NLU and DST) [50, 51, 52, 53] and word-level policy (coupling Policy and NLG). As shown in Figure 3, the evolutionary trajectory of each sub-module in TOD has been underpinned by the rapid progress of neural language models. Notably, a pivotal transformation has been observed in the transition from conventional neural language models (NLMs) to the more versatile pre-trained language models (PLMs). This shift has significantly contributed to the enhancement of the dialogue system’s capabilities and performance.
@@
Moreover, there are two obvious trends in this progress. First, TOD initially is designed to perform specific tasks within a predefined domain, however, as the PLMs emerged and gained traction, this paradigm shifted towards pre-train and then finetune [26]. By taking advantage of intricate patterns and world knowledge in the parameters of PLMS, the following method based on PLMs can generalize across tasks [21] and domains [41], even languages [147]. As tasks increase in complexity, they give rise to emerging challenges, including but not limited to insufficient availability of annotated data and the exacerbation of error propagation phenomena. To address these issues, inspired by the notable achievements of PLMs, a considerable body of research endeavors to employ an akin approach to conduct pre-training or in-domain finetuning on corpora specific to these sub-tasks in TOD simultaneously, constituting the second discernible trend. For example, SimpleTOD [9] uses a single, causal language model trained on all sub-tasks recast as a single sequence prediction problem, leading to better performance at all sub-tasks and robustness to noises. Consequently, these obstacles between distinct components and tasks gradually diminish, and more unified end-to-end TODs are on the way (§ 4.1).
3.2 The Development of ODD
@@
Distinguishing from TOD, open-domain dialogue systems (ODDs) aim to perform chit-chat with users without task and domain restrictions and are usually fully data-driven in an end2end manner [136, 1]. In 2016, soon after the emergence of the seq2seq framework, the Alexa Price challenge was proposed to advance the development of ODD that is able to converse coherently and engagingly with users on open-end topics such as sports, politics, and entertainment [148]. Generally, there are three types of open-domain dialogue systems: generative dialogue systems [75, 78, 81, 93, 149], retrieval-based dialogue systems [95, 98, 104, 107, 111], and hybrid dialogue systems [112, 113, 120]. In detail, the retrieval-based systems retrieve predetermined responses from a fixed set in response to user input. In contrast, generative systems construct responses from scratch using language models, resulting in diverse and contextually appropriate outputs. Hybrid systems combine both approaches, using retrieval for efficiency and generative methods for flexibility, aiming to achieve a balance between accuracy and creativity. The choice of system type depends on factors such as the application’s requirements, available resources, and the desired user experience. Some representative works are summarized in Figure 3.
@@
Similar to the TOD, ODD also went through a shift from NLM to PLM since the appearance of GPT [11], which operates in a decoder-only framework and is trained by predicting the next token auto-regressively. Most ODD based on NLM utilizes two typical recurrent neural networks: LSTM [19] and GRU [20], such as response selection tasks [111], knowledge-grounded dialogue [76, 83, 150], persona-consistent dialogue [78, 80, 79, 151], keywords [82] and topic [81] guided dialogue. A similar trend continues except the backbone of the dialogue system becomes PLMs after the appearance of BERT [24] and GPT2 [25]. Inspired by the exceptional capability learned during the pre-training stage, lots of works either directly pre-train the model from scratch using the dialogue dataset or further conduct in-domain finetuning using the current PLM as the initial checkpoint. The model after pretraining on tremendous dialogue datasets is formally defined as pre-trained dialogue model (PDM).
@@
Distinguish from TOD which usually adopts encoder-only PLM, there are usually two specific architectures of different PDMs: encoder-decoder [30, 85, 29] and decoder-only [13, 94], especially for generative and hybrid ODD. The distinction here can be attributed to the core nature and objectives of TOD and ODD, as the components of TOD primarily revolve around natural language understanding tasks such as NLU, DST, and DPL, while ODD focuses on natural language generation tasks. Specifically, as shown in the bottom part of Figure 3, Meena [30] from Google, Blenderbot [29] from Meta, and EVA [85] from Tsinghua University are all based on encoder-decoder architecture, which demonstrates strong understanding and generation capability simultaneously. In addition, most decoder-only PDMs follow the architecture of GPT2 [25] such as DialoGPT [13] and CDialoGPT [94] for English and Chinese dialogue respectively. Back to this moment, it is observed that most industrial and academic institutions choose to select decoder-only architecture aiming to improve since the scaling law has been widely observed in causal decoders [152].
As the size of the pre-training corpus and the model itself continue to expand, researchers have observed a concurrent improvement in the performance of language models. Throughout this evolution, certain studies have sought to integrate various types of dialogues using meticulously crafted pre-training objectives. The objective is to enable the resulting model to attain enhanced performance in both Task-oriented Dialogue (TOD) and Open-domain Dialogue (ODD) simultaneously [92, 153].
4 Fusion between LLM and DM
@@
Benefiting from scaling law [33], the performance of pre-trained language models (PLM) can be significantly improved by increasing the model size and training data volume. This scaling enables language models to achieve higher levels of understanding and generation capabilities simultaneously. For example, the BERT [24] and GPT series models [11] exemplify this capability, leading to three kinds of fusions during the scaling: 1): fusion within task-oriented dialogue system; 2) fusion between two types of dialogue systems: task-oriented dialogue system and open-domain dialogue system; and 3) fusion between large language model and dialogue model.
4.1 Fusion within TOD
As PLM demonstrates powerful capabilities at different downstream tasks, a lot of work starts to fuse several sub-tasks in TOD (e.g., NLU, DST, DPL, and NLG) and propose a unified end-to-end framework to solve them simultaneously [22, 154, 64, 155, 156, 157], starting from the earliest work NLM (e.g, LSTM, RNN) to the latter PLM (e.g, GPT2, T5), as shown in Table II. By observing Table II, we find that there are several trends accompanied by the shift of used language models: In general, since the appearance of PLM, there is an increasing interest and attention to the end-to-end task-oriented dialogue system as we observe more works appear. Most of them adopt encoder-decoder or decoder-only architectures by formulating each sub-task as a sequence-to-sequence task while leveraging both the understanding and generation capability of language models [158, 157].
@@ 4.1 Fusion within TOD
Specifically, at the task level in Table II, it is necessary to incorporate the final response generation task (a.k.a, NLG), but the other three sub-tasks are not essential for end-to-end TOD, given that there are multiple methods that do not involve any of the other three tasks. According to the frequency of each task in the different methods, the order relationship is NLG $>$ DST $>$ DPL $>$ NLU. Besides the NLG, DST which tracks the belief state during the conversation becomes more important, since the task-oriented dialogues include many interactions with a specific goal. Furthermore, apart from four key sub-tasks in traditional TOD, we can find more and more additional tasks are optimized with the mentioned key sub-tasks simultaneously. For example, Kulhanek et al. [159] use BC, SAC, and IC to further improve the diversity and quality of responses. Similarly, He et al. [157] use MLM and SRM to learn from large-scale dialog corpora with limited annotations. On the other hand, at the data level, SOLOIST [64], as the first work to investigate the effects of pre-training on heterogeneous dialog corpora, demonstrates great adaptability to accomplish new tasks efficiently. Afterward, lots of studies utilize additional data at the pre-training stage [64, 159, 157] or fine-tuning stage [155, 156], such as He et al. [156] propose GALAXY to explicitly learn dialog policy from limited labeled dialogs and large-scale unlabeled dialog corpora via semi-supervised learning. The only exception is OPERA [160], which proposes a novel dataset OB-MultiWOZ by incorporating question-answering tasks into task-oriented dialogues.
@@
TABLE II: An overview of common E2E Task-oriented Dialogue System, listed in time order.
Model Year Base Model NLU DST DPL NLG Add Tasks* Add Data Evaluation Dataset
NN [22] 2017 LSTM, RNN, CNN Wizard-of-Oz Data
MOSS [154] 2019 GRU CamRest676, LaptopNetwork
MinTL [161] 2020 T5, BART MultiWOZ 2.0
SimpleTOD [162] 2020 GPT2 MultiWOZ 2.0
SOLOIST [64] 2021 GPT2 Cont CamRest676, MultiWOZ 2.0
UBAR [163] 2021 DistilGPT2 MultiWOZ 2.0, 2.1
MTTOD [164] 2021 T5 SP MultiWOZ 2.0, 2.1
AuGPT [159] 2021 GPT2 BC, SAC, IC MultiWOZ 2.0, 2.1
GALAXY [156] 2022 UniLM RS, CR In-Car, MultiWOZ 2.0, 2.1
PPTOD [155] 2022 T5 MultiWOZ 2.0 2.1
OPERA [160] 2022 T5 QA OB-MultiWOZ
SPACE-3 [157] 2022 UniLM MLM,SRM MultiWOZ, CamRest676, In-Car
* Additional task objective abbreviations: Cont-Contrastive Loss, SP–Span Prediction, BC–Belief Consistency, SAC–System Action Classification)
IC–Intent Classification, RS–Response Selection, CR–Consistency Regularization, QA–Question Answering, MLM-Masked Language Modeling
SRM-Semantic Region Modeling
@@
The progression of task-oriented end-to-end dialogue systems has been characterized by notable developments in several key areas. As dialogue systems evolved, encompassing the data and tasks, there emerged a growing emphasis on their evaluation in settings that better simulate real-world scenarios. There are some works that recognize the inherent challenge of realistic settings, such as UBAR [163] is evaluated as a fully end-to-end task-oriented dialogue system where its dialog context has access to user utterances and all content it generated such as belief states, system acts, and system responses. Nonetheless, many studies directly assess the performance of different methods either employing various backbones or incorporating additional data [64, 155, 159]. Given that the choice of pre-trained language models (PLMs) and the pre-training data significantly impact the final performance, the comparison is not that fair, which underscored the necessity for benchmarking in more realistic and fair settings.
The evolution of task-oriented end-to-end dialogue systems has been marked by significant advancements in data, the unification of different tasks, a transition to more realistic settings, and sequence-to-sequence modeling. As the understanding and generation capability of language models keep improving with the scaling of data size and model size, the future holds promise for more versatile and effective task-oriented dialogue systems that can understand and respond to user needs in a wide range of domains and languages [139, 165]. More details can be found in the corresponding survey paper [166].
4.2 Fusion between TOD and ODD

### v2

@@ 3. The Independent Development of TOD & ODD
Figure 2. Taxonomy of LM-based Dialogue System, with representative LMs/Method of each stage: 1) Independent development of TOD; 2) Independent development of ODD; and 3) Fusion between TOD and ODD.
3.1. The Development of TOD
@@
In the nascent stages of task-oriented dialogue system development, exemplified by systems like GUS (Bobrow et al., 1977) and ALICE (Wallace, 2009), the technology relied heavily on rule-based approaches. These early systems operated within single domains and adhered to pre-defined decision trees or rule sets. For example, GUS was constrained to specific domains such as flight information retrieval (Bobrow et al., 1977), while ALICE was designed to simulate natural language conversation, following a script-like structure based on pattern matching and predefined responses (Wallace, 2009). These systems showcase impressive capabilities within their limited contexts but struggle to handle nuanced or complex conversations. They often require extensive manual crafting of rules and lack the ability to generalize beyond their narrow domains, making them ill-equipped for real-world applications with varying user intents and open-ended dialogues.
@@
With the advent of the neural language model (NLM) and subsequent advancements in recurrent neural networks (RNNs) (Hochreiter and Schmidhuber, 1997; Cho and et al., 2014), the landscape of task-oriented dialogue systems underwent a first evolution. The introduction of RNNs allows for more flexible and adaptive responses, enabling systems to process and generate natural language in a contextually aware manner. Dialogue systems now exhibit a broader spectrum of capabilities, embracing multi-domain functionality and accommodating diverse user intents. Unlike their predecessors, these modern systems can handle complex, intertwined tasks within a single conversation. For instance, a task-oriented dialogue system might seamlessly assist a user in booking a flight, reserving a hotel room, and planning an itinerary, all within the same interaction. This expanded capability has been made possible by the integration of four different components, forming a common pipeline for existing task-oriented dialogue systems: Natural Language Understanding (NLU) (Qin et al., 2021), Dialogue State Tracking (DST) (Jacqmin and et al., 2022), Dialogue Policy Learning (DPL) (Kwan and et al., 2023), and Natural Language Generation (NLG) (Li and Yao, 2021)  [^5] https://github.com/yizhen20133868/Awesome-TOD-NLG-Survey.git. Nevertheless, it is important to note that most studies at this stage focus on individual components and propose task-specific designs, leading to limited scalability and poor generalization. Moreover, there are some works try to combine consecutive components, such as word-level DST (coupling NLU and DST) (Henderson et al., 2014b; Mrkšić et al., 2017, 2015; Perez and Liu, 2017) and word-level policy (coupling Policy and NLG). As shown in Figure 2, the evolutionary trajectory of each sub-module in TOD has been underpinned by the rapid progress of neural language models.
@@
Notably, a pivotal transformation has been observed in the transition from conventional neural language models (NLMs) to the more versatile pre-trained language models (PLMs). This shift has significantly contributed to the enhancement of the dialogue system’s capabilities and performance. Specifically, there are two obvious trends in this progress. First, TOD initially is designed to perform specific tasks within a predefined domain, however, as the PLMs emerged and gained traction, this paradigm shifted towards pre-train and then finetune (Liu and et al., 2023a). By taking advantage of intricate patterns and world knowledge in the parameters of PLMS, the following method based on PLMs can generalize across tasks (Peng and et al., 2017) and domains (Hakkani-Tür and et al., 2016), even languages (Wang and et al., 2021a). As tasks increase in complexity, they give rise to emerging challenges, including but not limited to insufficient availability of annotated data and the exacerbation of error propagation phenomena. To address these issues, inspired by the notable achievements of PLMs, a considerable body of research endeavors to employ an akin approach to conduct pre-training or in-domain finetuning on corpora specific to these sub-tasks in TOD simultaneously, constituting the second discernible trend. For example, SimpleTOD (Hosseini-Asl et al., 2020) uses a single, causal language model trained on all sub-tasks recast as a single sequence prediction problem, leading to better performance at all sub-tasks and robustness to noises. Consequently, these obstacles between distinct components and tasks gradually diminish, and more unified end-to-end TODs are on the way (§ 4.1).
3.2. The Development of ODD
@@
Distinguishing from TOD, open-domain dialogue systems (ODDs) aim to perform chit-chat with users without task and domain restrictions and are usually fully data-driven in an end2end manner (Chen and et al., 2017; Huang and et al., 2020). Generally, there are three types of open-domain dialogue systems: generative dialogue systems (Li et al., 2016; Zhang et al., 2020a; Wang and et al., 2022), retrieval-based dialogue systems (Wu and et al., 2017; Tan et al., 2016; Henderson and et al., 2020; Whang and et al., 2021; Gao et al., 2020), and hybrid dialogue systems (Zhou et al., 2020; Zhang and et al., 2019; Lin et al., 2021). In detail, the retrieval-based systems retrieve predetermined responses from a fixed set in response to user input. In contrast, generative systems construct responses from scratch using language models, resulting in diverse and contextually appropriate outputs. Hybrid systems combine both approaches, using retrieval for efficiency and generative methods for flexibility, aiming to achieve a balance between accuracy and creativity. The choice of system type depends on factors such as the application’s requirements, available resources, and the desired user experience. Some representative works are summarized in Figure 2.
@@
As early as 2016, shortly after the introduction of the seq2seq framework, the Alexa Prize challenge was launched to promote the development of open-domain dialogue (ODD) systems capable of engaging in coherent and engaging conversations with users on open-ended topics like sports, politics, and entertainment (Chen and et al., 2018). Since then, ODD also went through a shift from NLM to PLM since the appearance of GPT (Radford et al., 2018), which operates in a decoder-only framework and is trained by predicting the next token auto-regressively. During the era of NLM, most ODD utilizes two typical recurrent neural networks: LSTM (Hochreiter and Schmidhuber, 1997) and GRU (Cho and et al., 2014) such as response selection tasks (Gao et al., 2020), knowledge-grounded dialogue (Parthasarathi and Pineau, 2018; Ghazvininejad et al., 2018; Wang et al., 2023a), persona-consistent dialogue (Li et al., 2016; Kottur et al., 2017; Zhang et al., 2018; Chen and et al., 2023), keywords (Zhong et al., 2021) and topic (Zhang et al., 2020a) guided dialogue. A similar trend continues except the backbone of the dialogue system becomes PLMs after the appearance of BERT (Devlin and et al., 2019) and GPT2 (Radford et al., 2019). Inspired by the exceptional capability learned during the pre-training stage, lots of works either directly pre-train the model from scratch using the dialogue dataset or further conduct in-domain finetuning using the current PLM as the initial checkpoint. The model after pretraining on tremendous dialogue datasets is formally defined as pre-trained dialogue model (DM).
@@
There are two common architectures for different PDMs: encoder-decoder (Adiwardana et al., 2020; Gu and et al., 2022; Roller et al., 2021) and decoder-only (Zhang et al., 2020b; Wang et al., 2020a), particularly for generative and hybrid ODD systems. This contrasts with TOD systems, which typically use encoder-only PLMs. The distinction here can be attributed to the core nature and objectives of TOD and ODD, as the components of TOD primarily revolve around natural language understanding tasks such as NLU, DST, and DPL, while ODD focuses on natural language generation tasks. Specifically, as shown in the bottom part of Figure 2, Meena (Adiwardana et al., 2020) from Google, Blenderbot (Roller et al., 2021) from Meta, and EVA (Gu and et al., 2022) from Tsinghua University are all based on encoder-decoder architecture, which demonstrates strong understanding and generation capability simultaneously. In addition, most decoder-only PDMs follow the architecture of GPT2 (Radford et al., 2019) such as DialoGPT (Zhang et al., 2020b) and CDialoGPT (Wang et al., 2020a) for English and Chinese dialogue respectively. Back to this moment, it is observed that most industrial and academic institutions choose to select decoder-only architecture aiming to improve since the scaling law has been widely observed in causal decoders.
3.3. Summary
At this stage, TOD and ODD exhibit distinct boundaries in the data, task and model architectural designs. While TOD focuses on goal-driven interactions (e.g., hotel bookings) with structured pipelines, ODD prioritizes free-form, contextually rich conversations. Advances in pre-trained language models (PLMs) have propelled both paradigms toward more versatile and capable conversational agents. However, their development remains characterized by fragmented methodologies and diverse experimentation, with significant exploration and uncertainty in key areas: i) data: Domain-specific and task-specific corpora dominate TOD, whereas ODD relies on open-domain datasets, results in significant differences in data schema and format between the two; ii) task: TOD modularizes sub-tasks (NLU, DST, NLG), while ODD emphasizes end-to-end generation; and iii) backbone model: most of sub-tasks in TOD adopt encoder-only architectures (e.g., BERT) while ODD favors decoder-only models (e.g., GPT).
Due to the existing complexity and uncertainty, there is no one-size-fits-all solution for TOD and ODD. Fortunately, the advent of PLMs has opened a promising path toward more unified, scalable, and powerful dialogue systems. The development of dialogue systems afterwards is, therefore, an ongoing process aimed at reducing this complexity and uncertainty, ultimately leading to a universal approach that evolves through a bottom-up integration.
4. Fusion of Dialogue Systems
@@
Benefiting from scaling law (Kaplan et al., 2020), the performance of pre-trained language models (PLM) can be significantly improved by increasing the model size and training data volume. This scaling enables language models to achieve higher levels of understanding and generation capabilities simultaneously. For example, the BERT (Devlin and et al., 2019) and GPT series models (Radford et al., 2018) exemplify this capability, leading to three kinds of fusions during the scaling: 1): fusion within task-oriented dialogue system; 2) fusion between two types of dialogue systems: task-oriented dialogue system and open-domain dialogue system; and 3) fusion between large language model and dialogue model.
4.1. Fusion within TOD
Such modular TOD faces several critical limitations, including cascading errors from different components (e.g., NLU misclassifications corrupting the later DST) and high costs for domain-specific adaptation, resulting in low scalability and generalizability. The advert of seq-to-seq framework and especially the appearance of PLM make it possible to unify all sub-tasks into shared architectures, reducing extensive manual feature engineering for each module in traditional TOD. Due to extensive world knowledge pre-trained in PLM, it brings much better scalability, generalizability and also the overall performance.
Specifically, as PLMs demonstrates exceptional performance across various natural language processing downstream tasks, an increasing number of studies seek to leverage them to address multiple sub-tasks in TOD (e.g., NLU, DST, DPL, and NLG) simultaneously within a unified end-to-end framework (Wen et al., 2017b; Liang et al., 2020; Peng et al., 2021; Su et al., 2022; He and et al., 2022a, b), starting from the earliest work NLM (e.g, LSTM, RNN) to the latter PLM (e.g, GPT2, T5), as shown in Table 3. Several conclusion can be drawn from the results. In general, there is a significant increase in the number of studies aiming to solve multiple sub-tasks in TOD simultaneously (Peng et al., 2021; Hosseini-Asl et al., 2020; Li and et al., 2023). Some also explore incorporating auxiliary tasks, such as masked language modeling (MLM) (He and et al., 2022b) and response selection (RS) (He and et al., 2022a), to enhance overall performance. All of these approaches benefit from the seq-to-seq framework, leveraging various base models based on encoder-decoder or decoder-only architectures.
@@ 4.1. Fusion within TOD
In detail, the inherent characteristics of each task and the diverse corresponding data formats necessitate distinct strategies for effectively fusing multiple tasks in end-to-end TOD frameworks. On the task level, it is necessary to incorporate the final response generation task (a.k.a, NLG), but the other three sub-tasks are not essential for end-to-end TOD, given that there are multiple methods that do not involve any of the other three tasks. According to the frequency of each task in the different methods, the order relationship is NLG $>$ DST $>$ DPL $>$ NLU. Besides the NLG, DST which tracks the belief state during the conversation becomes more important, since the task-oriented dialogues include many interactions with a specific goal. Furthermore, apart from four key sub-tasks in traditional TOD, we can find more and more additional tasks are optimized with the mentioned key sub-tasks simultaneously. For example, Kulhanek et al. (Kulhánek and et al., 2021) use BC, SAC, and IC to further improve the diversity and quality of responses. Similarly, He et al. (He and et al., 2022b) use MLM and SRM to learn from large-scale dialog corpora with limited annotations. On the the data level, SOLOIST (Peng et al., 2021), as the first work to investigate the effects of pre-training on heterogeneous dialogue corpora, demonstrates great adaptability to accomplish new tasks efficiently. Afterward, lots of studies utilize additional data at the pre-training stage (Peng et al., 2021; Kulhánek and et al., 2021; He and et al., 2022b) or fine-tuning stage (Su et al., 2022; He and et al., 2022a), such as He et al. (He and et al., 2022a) propose GALAXY to explicitly learn dialog policy from limited labeled dialogs and large-scale unlabeled dialog corpora via semi-supervised learning. The only exception is OPERA (Li and et al., 2023), which proposes a novel dataset OB-MultiWOZ by incorporating question-answering tasks into task-oriented dialogues.
@@
Table 3. An overview of common E2E Task-oriented Dialogue System, listed in time order.
Model Year Base Model NLU DST DPL NLG Add Tasks* Add Data Evaluation Dataset NN (Wen et al., 2017b) 2017 LSTM, RNN, CNN Wizard-of-Oz Data MOSS (Liang et al., 2020) 2019 GRU CamRest676, LaptopNetwork MinTL (Lin et al., 2020b) 2020 T5, BART MultiWOZ 2.0 SimpleTOD (Hosseini-Asl et al., 2020) 2020 GPT2 MultiWOZ 2.0 SOLOIST (Peng et al., 2021) 2021 GPT2 Cont CamRest676, MultiWOZ 2.0 UBAR (Yang et al., 2021) 2021 DistilGPT2 MultiWOZ 2.0, 2.1 MTTOD (Lee, 2021) 2021 T5 SP MultiWOZ 2.0, 2.1 AuGPT (Kulhánek and et al., 2021) 2021 GPT2 BC, SAC, IC MultiWOZ 2.0, 2.1 GALAXY (He and et al., 2022a) 2022 UniLM RS, CR In-Car, MultiWOZ 2.0, 2.1 PPTOD (Su et al., 2022) 2022 T5 MultiWOZ 2.0 2.1 OPERA (Li and et al., 2023) 2022 T5 QA OB-MultiWOZ SPACE-3 (He and et al., 2022b) 2022 UniLM MLM,SRM MultiWOZ, CamRest676, In-Car * Additional task objective abbreviations: Cont-Contrastive Loss, SP–Span Prediction, BC–Belief Consistency, SAC–System Action Classification) IC–Intent Classification, RS–Response Selection, CR–Consistency Regularization, QA–Question Answering, MLM-Masked Language Modeling SRM-Semantic Region Modeling
@@
Building on these task- and data-level fusions, the evolution of task-oriented end-to-end dialogue systems also emphasizes more realistic evaluation scenarios. There are some works that recognize the inherent challenge of realistic settings, such as UBAR (Yang et al., 2021) is evaluated as a fully end-to-end task-oriented dialogue system where its dialog context has access to user utterances and all content it generated such as belief states, system acts, and system responses. Nonetheless, many studies directly assess the performance of different methods either employing various backbones or incorporating additional data (Peng et al., 2021; Su et al., 2022; Kulhánek and et al., 2021). Given that the choice of pre-trained language models (PLMs) and the pre-training data significantly impact the final performance, the comparison is not that fair, which underscored the necessity for benchmarking in more realistic and fair settings.
The evolution of task-oriented end-to-end dialogue systems has been marked by significant advancements in data, task and evaluations, which offers notable advantages and also potential limitations. On the positive side, unifying sub-tasks such as NLU, DST, DPL and NLG into a single model reduces cascading errors found in modular pipelines by enabling joint optimization and mitigating the propagation of misalignments between components. Moreover, the unified training format enhances scalability and generalizability, allowing these systems to adapt more effectively to unseen tasks and domains. In contrast, there are still some limitations in terms of data curation and evaluation. To train the model in an end-to-end manner on diverse tasks, it is necessary to maintain a unified data format and pool, leading to increased costs in data curation and annotation. Besides that, evaluating and optimizing such integrated systems becomes more complex since it is difficult to optimize each sub-module independently. Nevertheless, as the understanding and generation capability of language models keep improving with the scaling of data size and model size, the future holds promise for more versatile and effective task-oriented dialogue systems that can understand and respond to user needs in a wide range of domains and languages (Kwan and et al., 2023; Ni et al., 2023). More details can be found in the corresponding survey paper (Qin and et al., 2023).
4.2. Fusion between TOD and ODD
The emergence of end-to-end task-oriented dialogue (TOD) systems has catalyzed the development of unified dialogue frameworks capable of integrating both task-oriented and open-domain dialogue (ODD) functionalities. This architectural convergence — rooted in the adoption of sequence-to-sequence (seq-to-seq) models — addresses critical limitations of traditional systems while enabling holistic user experiences. Historically, TOD and ODD systems operated in isolation: TOD focused on task completion (e.g., booking flights), while ODD prioritized social interaction (e.g., chit-chat). However, real-world conversations inherently blend these modalities, such as when users discuss travel preferences while completing a booking. Discrete architectures struggled to manage such interleaved interactions, resulting in fragmented user experiences and prohibitive computational costs due to duplicated infrastructure.